A stone is dropped from the roof of a high building. A second stone is dropped 1.42s later.

How far apart are the stones when the second one has reached a speed of 14.3 m/s?

d=???m
Answer #1

Final velocity of second stone = 14.3 m/s

v = u + at

v = at

t = 14.3 / 9.8

t = 1.4592 sec

so,

d = 0.5 * 9.8 * 1.45922

d = 10.433 m

so, Total time for which first stone has been moving = 1.4592 + 1.42 = 2.8792 sec

so,

distance covered = 0.5 * 9.8 * 2.87922

distance covered = 40.619 m

so,

stones are apart by

d = 40.619 - 10.433

d = 30.186 m